OpenStudy (anonymous):

square root of 49x^2 y^10

OpenStudy (anonymous):

\[\LARGE \sqrt{49x^2y^{10}}=\sqrt{49}\cdot \sqrt{x^2}\cdot \sqrt{y^{8}\cdot y^2}=...\]

OpenStudy (anonymous):

can you do it now ? :)

OpenStudy (anonymous):

@StephanieEatsPants do you understand it? :) ... if you don't, just let me know.

OpenStudy (anonymous):

... ok let's assume you don't speak english, so I have to do it lol :) \[\Large \sqrt{49}\cdot \sqrt{x^2}\cdot \sqrt{y^8\cdot y^2}=7\cdot x\cdot y^4\cdot y\] actually there was unnecessary to split y^10 but I didn't notice that. so the answer is 7xy^5